OpenStudy (anonymous):

write the decimal as a fraction in simplest form 0.52

OpenStudy (anonymous):

\[\frac{0.52}{1}*\frac{100}{100}=\frac{52}{100}=\frac{13}{25}\]

OpenStudy (anonymous):

13/25

OpenStudy (anonymous):

13/25